# Multi-Level Querying using A Knowledge Pyramid

### v2

@@ 3.2 Knowledge Completion
where $n$ is the dimension of the space, $\mathbf{X}$ denotes a phrase embedding, and $\mathbf{\mu}$ and $\mathbf{\Sigma}$ are the mean and covariance matrix, respectively. The semantic distributions of the Ontology and knowledge graph are then modeled as $\mathbf{X}_{\mathcal{O}}\sim\mathcal{N}(\mathbf{\mu}_{\mathcal{O}},\mathbf{\Sigma}_{\mathcal{O}})$ and $\mathbf{X}_{\mathcal{K}}\sim\mathcal{N}(\mathbf{\mu}_{\mathcal{K}},\mathbf{\Sigma}_{\mathcal{K}})$, respectively. The Kullback–Leibler (KL) divergence can then be measured on each given phrase embedding $\mathbf{X}$ as
@@
At a specific location $\mathbf{X}$, we define a function to rate the priority to extract entities from the knowledge graph to the Ontology as
@@
The rating outcomes are depicted in Figure 2, illustrating that this function effectively identifies areas where the knowledge graph contains dense information that is comparatively lacking in the Ontology. We then apply k-medoids clustering on knowledge graph triplets within these area to locate new Ontology class and attributes that can be integrated into the Ontology schema, along with corresponding instances. This process can be performed iteratively, allowing for continuous refinement and improvement. As a result of these interactions between the two layers, the comprehensiveness of the pyramid is progressively enhanced.
3.3 Knowledge Condensation
The process of Knowledge Completion aims to enhance the richness of information in a bottom-up manner. In addition, we perform Knowledge Condensation to improve the compactness of the pyramid in a top-down approach. The principle is to leverage the well-structured knowledge in the higher layers to eliminate redundant information present in the lower layers. This is implemented by using each class-attribute in the Ontology layer as an anchor and explore its neighborhood to identify a set of $k$ nearest knowledge graph triplets $\{\mathbf{X_{k}}\}$. A prompt function is then applied to instruct the LLM to summarize them into compact ones as
@@ 3.3 Knowledge Condensation
By repeating this process for all anchors, we obtain the condensed knowledge graph layer, as depicted in Figure 2. The original triplets such as “Prof. A works in Z” and “Prof. A is with Z institue” have been merged into more compact form as a relation of “is_a_member_of”. It is evident that the compactness of the pyramid has been significantly improved.
4 Multi-Level Querying
@@
1:function RetrievePolyRAG($Q$)
2:     $qres\leftarrow Query_{onto}(Q,schema)$
3:     $C\leftarrow Onto(qres)$
4:     if  $C$ is empty then
5:         $triples\leftarrow Emb_{kg}(Q,\mathcal{K})$
6:         $agreement\leftarrow Query_{kg}(Q,triples)$
@@
10:              $C\leftarrow Emb_{rag}(Q,\mathcal{T})$
@@
4.1 Search at Ontology Layer
@@
At the Ontology Layer, we can utilize SparQL as the query language. SparQL is a well-defined language that provides precise results when queries are properly structured. To achieve this, we can guide the LLM to generate the SparQL query by using the following prompt function
@@
$Query_{onto}(Q,schema)$: Given a input question {$Q$} within the domain {domain}, please formulate a SPARQL query to retrieve the answer based on the provided Ontology schema. The namespace is {$schema.base$}, and the classes are {$schema.class$}. The object properties between classes are {$schema.op$}, and the classes may also have data properties such as {$schema.dp$}.
@@
For the example of a query question like “Who is currently working in CS Department and was graduated from Cambridge University?”, the LLM extracts the related attributes of “works_in” and “graduated_from” with the given $schema$, and generate a SPARQL query to search a staff who satisfies the relevant condition in $\mathcal{O}$. The executing results may include a list of names, or simply empty due to either the lack of knowledge in Ontology layer, or the mistake of providing unavailable SPARQL due to the limitation of LLM itself.
4.2 Search at Knowledge Graph Layer
@@
At the knowledge graph layer, we employ a retrieval approach similar to the embedding-based retrieval utilized in NaiveRAG. However, instead of using chunks, we work with triplets. Once the matching triplets are retrieved, we utilize a prompt function to assess the Language Model’s agreement on whether the question has been answered adequately as
@@
$Query_{kg}(Q,triples)$: Given a question {$Q$} and the context information provided by the matched triples from the knowledge graph {$triples$}, please justify whether the provided information is sufficient to accurately answer the question. Respond with either “Yes” or “No” to provide your justification.
@@
For example, given a query of “which CS staff has interest in cloud computing?”, the LLM may justify a set of triples as “agree” if they include information like “Prof. A works in CS Department, Prof. A published on cloud computing journal, … etc.”; inversely, the justification might be “disagree” if the triples do not provide the key knowledge related to this query, thus the searching will process to the next layer for a more relevant context in order to answer the query.
@@
Table 1: Performance comparison with SOTA methods on the AcadChall and R-FLUE-FiQA datasets.
@@
5 Experiments
5.1 Experimental Setup
@@
Our experimentation involves two distinct domain-specific benchmarks. The first benchmark is Academia Challenge (AcadChall), which has been built by ourselves and focuses on the academic domain. It encompasses a comprehensive collection of data obtained from XXX university, including information about 31 departments, 1,319 faculty members, and 2,061 courses. Specifically, AcadChall consists of 512 MCQ and MAQ questions that cover topics related to teaching and research. These questions are intentionally designed to be more challenging compared to existing benchmarks, as each question presents eight answer choices. This design aims to provide a more rigorous assessment of the models’ precision. The second benchmark is FiQA, which is widely recognized and utilized in the finance domain. Our experimentation follows the dataset split provided by FLUE benchmark Shah et al. (2022), resulting in the creation of the R-FLUE-FiQA dataset comprising 1,705 open-ended questions. To enhance knowledge completion and condensation, we employ the methods proposed in Section 3 to construct Ontologies, knowledge graphs, and pyramids based on the two datasets. For more detailed information about the two benchmarks, please refer to the Appendix B.
@@
We employ instructor-xl Su et al. (2022) as our vector base for embedding search and the cosine similarity is used as the metric Wilkinson and Hingston (1991). The $k$ values for KG layer and raw texts layer are empirically set to 10 and 5, respectively.
Table 2: The impact of Knowledge Completion (CPL) and Condensation (CND) by F1 scores.
Method Vicuna-7B Gemini-Pro GPT-3.5 GPT-4
Naive RAG 0.4673 0.4214 0.5315 0.4651
Baseline 0.5734 0.6899 0.7905 0.7884
$+$ CPL 0.5874 0.7869 0.7939 0.7763
$+$ CND 0.5766 0.7439 0.7943 0.7873
$+$ CPL $+$ CND 0.5967 0.8107 0.8039 0.8109
5.2 Evaluation Metrics
@@
Our evaluation process encompasses several metrics to assess the performance, including: 1) For MCQs and MAQs, we employ metrics such as Precision, Recall, and F1 score. 2) For open-ended questions, we utilize BLEU and BERT similarity, calculated based on MiniLM embeddings Wang et al. (2020a). 3) We introduce a novel metric called HitRate ($HR$), which quantifies the proportion of correct entities present in the response. These metrics collectively allow for a comprehensive assessment of the system’s effectiveness across different question types and response formats.
Table 3: The influence of different knowledge layers on the AcadChall benchmark.
Pyramid Layer Vicuna-7b Gemini-Pro GPT-3.5 GPT-4
Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1
Ontology 0.4383 0.6242 0.4445 0.3386 0.3089 0.3157 0.4107 0.3945 0.3979 0.3500 0.4369 0.3677
KG 0.4365 0.8581 0.5133 0.6388 0.4591 0.5099 0.7326 0.6340 0.6433 0.6327 0.7065 0.6238
Raw Texts 0.3621 0.8386 0.4673 0.5152 0.3968 0.4214 0.5105 0.5791 0.5315 0.5032 0.5094 0.4651
Ontology + Raw Texts 0.4535 0.8393 0.5368 0.5905 0.4662 0.5013 0.6890 0.5818 0.6020 0.6464 0.6779 0.6303
KG + Raw Texts 0.3673 0.8758 0.4840 0.6890 0.5428 0.5821 0.8339 0.6622 0.7071 0.8187 0.7729 0.7636
Ontology + KG 0.5251 0.8266 0.5608 0.7755 0.6697 0.6709 0.8567 0.7710 0.7939 0.6083 0.9032 0.6871
PolyRAG 0.5306 0.8337 0.5967 0.8372 0.8419 0.8107 0.8667 0.7860 0.8039 0.8711 0.7992 0.8109
5.3 Comparison to SOTA Methods

### v3

@@ III-B Knowledge Completion
where $n$ is the dimension of the space, $\mathbf{X}$ denotes a phrase embedding, and $\mathbf{\mu}$ and $\mathbf{\Sigma}$ are the mean and covariance matrix, respectively. The semantic distributions of the Ontology and knowledge graph are modeled as $\mathbf{X}_{\mathcal{O}}\sim\mathcal{N}(\mathbf{\mu}_{\mathcal{O}},\mathbf{\Sigma}_{\mathcal{O}})$ and $\mathbf{X}_{\mathcal{K}}\sim\mathcal{N}(\mathbf{\mu}_{\mathcal{K}},\mathbf{\Sigma}_{\mathcal{K}})$. The Kullback–Leibler (KL) divergence is then measured on each phrase embedding as
@@
At a location $\mathbf{X}$, we define a function to rate the priority to extract entities from the knowledge graph to the Ontology as
@@
The rating outcomes are depicted in Figure 4, illustrating that this function effectively identifies areas where the knowledge graph contains dense information that is comparatively lacking in Ontology. We then apply k-medoids clustering on knowledge graph triplets within this area to locate new Ontology classes and attributes that can be integrated into the Ontology schema, along with corresponding instances. This process can be performed iteratively, allowing for continuous refinement and improvement. As a result of these interactions between the two layers, the comprehensiveness of the pyramid is progressively enhanced. Figure 3 provides a partial view of the Ontology to illustrate the completion process. In addition to the previously mentioned modifications to affiliations, entity Class and their connections to Staff and Research Interests have also been identified. The inclusion of them are useful for student queries seeking advisors.
Figure 5: Illustration of multi-level querying: the two queries are addressed at the Ontology and Knowledge Graph layers, respectively, bypassing the lower levels.
III-C Knowledge Condensation
The process of Knowledge Completion is designed to enhance the richness and comprehensiveness of information within the knowledge pyramid in a bottom-up manner, which improves the overall “recall” when retrieving information using the pyramid framework. By enriching the lower layers, such as raw texts and knowledge graphs, we ensure that more relevant information is available for retrieval. In addition, we perform Knowledge Condensation to improve the compactness and efficiency of the pyramid in a top-down manner, thereby enhancing “precision” in information retrieval.
The principle behind this approach is to leverage the well-structured knowledge in the higher layers, particularly the Ontology layer, to eliminate redundant or less relevant information present in the lower layers. This is implemented by using each class attribute in the Ontology as an anchor and exploring its neighborhood within the knowledge graph to identify a set of the $k$ nearest knowledge graph triplets ${\mathbf{X_{k}}}$. By focusing on these closely related triplets, we can filter out extraneous data and refine the knowledge base to include only the most pertinent information. This top-down refinement ensures that the system provides precise answers by relying on well-defined concepts and relationships, ultimately improving the quality and relevance of the retrieved information. A prompt function is applied to instruct the LLM to summarize those triplets into compact ones as
@@ III-C Knowledge Condensation
By repeating this condensation process for all anchors in the Ontology layer, we systematically traverse and refine the corresponding neighborhoods in the knowledge graph. This iterative application results in the formation of the condensed knowledge graph layer, as depicted in Figure 4. In this condensed layer, redundant and semantically similar triplets are merged based on their relations to the anchor concepts defined in the Ontology. For example, original triplets such as “Prof. A works in Z” and “Prof. A is with Z institute”, which may have been represented separately in the knowledge graph due to variations in phrasing, are consolidated into a single, more standardized relation of “is_a_member_of”. This condensation not only reduces redundancy but also enhances the semantic clarity of the knowledge representation by unifying similar relationships under a common relation. Such standardization is crucial for maintaining consistency across the knowledge graph and for facilitating more efficient query processing.
IV Multi-Level Querying
@@
1:function PolyRAG($Q$)
2:     $sparQ\leftarrow Query_{onto}(Q,schema)$
3:     $result\leftarrow Onto(sparQ)$
4:     if  $result$ is empty then
5:         $triplets\leftarrow Emb{Retrive}_{kg}(Q,\mathcal{K})$
6:         $agreement\leftarrow Query_{kg}(Q,triplets)$
@@
10:              $result\leftarrow Emb{Retrive}_{rawtext}(Q,\mathcal{T})$
@@
IV-A Search at Ontology Layer
@@
At the Ontology Layer, we can utilize SPARQL as the query language. SPARQL is a well-defined language that provides precise results when queries are properly structured [30]. To achieve this, we can guide the LLM to generate the SPARQL query by using the following prompt function
@@
$Query_{onto}(Q,schema)$: Given a question {$Q$} within the {domain} domain, please formulate a SPARQL query to retrieve the answer based on the provided Ontology schema. The namespace is {$schema.base$}, and the classes are {$schema.class$}. The object properties between classes are {$schema.op$}, and the classes may also have data properties such as {$schema.dp$}.
@@
For a query question like “What are the research interests of Prof. James Wang?”, the LLM extracts the related attributes of “research_interest” with the given $schema$, and generate a SPARQL query to search a staff who satisfies the relevant condition in $\mathcal{O}$. The executing results may include a list of interests or simply be empty due to the lack of knowledge in the Ontology layer.
Figure 5 presents an example of a query handled at the Ontology layer. The query “Which staff in CS department graduated from MIT” is asking for the knowledge of a professor with two restrictions: “current_affiliation: CS department” and “past_affiliation: MIT”. This input query is fed into the LLM with the template $Query_{onto}(Q,schema)$. The search process executes SPARQL language to the pre-built ontology database in OWL format. If there exist available results, namely the multi-level querying gets the satisfied context, the process terminates and fills the query results as the knowledge context for question-answering tasks. Due to the accuracy of the knowledge presented at the Ontology layer, the answers seldom fail as long as the knowledge needed is formulated in the ontology database and the LLMs generate the precise SPARQL language for querying. Moreover, since the retrieved results are direct knowledge, which is commonly short phrases of keywords and straightforward to what the query is searching for, it reduced the intelligence requirements of the LLMs, making PolyRAG an easier choice for small and middle-scale LLMs’ inference in terms of the context length.
IV-B Search at Knowledge Graph Layer
@@
At the knowledge graph layer, we employ a retrieval approach similar to the embedding-based retrieval utilized in NaiveRAG. However, instead of using chunks, we work with triplets. Treating the triplets as natural language knowledge instead of the graph is being proven to be useful for KG-enhanced LLM question-answering by [27]. Once the matching triplets are retrieved through the embedding search, we utilize a prompt function to assess the Language Model’s agreement on whether the question has been answered adequately as
@@
$Query_{kg}(Q,triplets)$: Given a question {$Q$} and the context information provided by the matched triplets from the knowledge graph {$triplets$}, please justify whether the provided information is sufficient to accurately answer the question. Respond with either “Yes” or “No” to provide your justification.
@@
For example, given a query of “Which CS staff has interest in cloud computing?”, the LLM may justify a set of triplets as “agree” if they include information like “Prof. A works in CS Department, Prof. A published on cloud computing journal, … etc.”; inversely, the justification might be “disagree” if the triplets do not provide the key knowledge related. The searching will then be processed to the next layer for a more relevant context in order to answer the query.
Figure 5 presents an example of a query handled at the Knowledge Graph layer. The query “What are the research interests of Alice” is asking for several research areas, and through the retrieval of the first layer, there is no hit due to the lack of knowledge in the ontology database. Then, the process comes to the second stage, which is to perform the embedding search of top-k similar triplets according to the semantics of the query. Given that a relevant group of triplets is being retrieved, the LLM justification module agrees these triplets should act as the background contexts. Therefore, the LLM is capable of generating the correct answers with the help of precise and concise retrieval.
TABLE I: Dataset statistics of the two benchmarks.
Dataset AcadChall R-FLUE-FiQA
Domain Academia Finance
#Raw texts 5,019 17,072
Ontologies
     #Instances 25,481 225,085
#KG Triples 27,920 68,737
#QA pairs 512 1,705
QA type MCQ&MAQ Open-ended
@@
TABLE II: Performance comparison with SOTA methods on the AcadChall and R-FLUE-FiQA datasets.
@@
V Experiments
V-A Experimental Setup
@@
Our experimentation involves two distinct domain-specific benchmarks. The first benchmark is Academia Challenge (AcadChall), which has been built by ourselves and focuses on the academic domain. It encompasses a comprehensive collection of data obtained from XXX university, including information about 31 departments, 1,319 faculty members, and 2,061 courses. Specifically, AcadChall consists of 512 MCQ and MAQ questions that cover topics related to teaching and research. These questions are intentionally designed to be more challenging compared to existing benchmarks, as each question presents eight answer choices. This design aims to provide a more rigorous assessment of the models’ precision. The second benchmark is FiQA, which is widely recognized and utilized in the finance domain. Our experimentation follows the dataset split provided by FLUE benchmark [32], resulting in the creation of the R-FLUE-FiQA dataset comprising 1,705 open-ended questions. To enhance knowledge completion and condensation, we employ the methods proposed in Knowledge Pyramid Construction Section to construct Ontologies, knowledge graphs, and pyramids based on the two datasets. The detailed information about the two benchmarks is shown in Table I.
@@
We employ instructor-xl [55] for embedding and the cosine similarity is used as the metric [56]. The $k$ values for KG layer and raw texts layer are empirically set to 10 and 5, respectively.
V-B Evaluation Metrics
@@
Our evaluation process encompasses several metrics to assess the performance, including: 1) For MCQs and MAQs, we employ metrics such as Precision, Recall, and F1 score. 2) For open-ended questions, we utilize BLEU and BERT similarity, calculated based on MiniLM embeddings [57]. 3) We introduce a novel metric called HitRate ($HR$), which quantifies the proportion of correct entities present in the response. These metrics collectively allow for a comprehensive assessment of the system’s effectiveness across different question types and response formats.
V-C Comparison to SOTA Methods